# The ECME Algorithm Using Factor Analysis for DOA Estimation in Nonuniform Noise

### v2

@@ 4.1 Algorithm Procedure
For problem (5), we do not adopt the complete data $(\mathbf{Y},\mathbf{X})$ with $\mathbf{X}=\big{[}\mathbf{x}(1)~\cdots~\mathbf{x}(L)\big{]}$ in [8, 17] and adopt the simpler complete data $(\mathbf{X},\mathbf{V})$ with $\mathbf{V}=\big{[}\mathbf{v}(1)~\cdots~\mathbf{v}(L)\big{]}$ [^3] Numerical results show that the ECME algorithm with $(\mathbf{X},\mathbf{V})$ yields the same convergence as that with $(\mathbf{Y},\mathbf{X})$. However, the ECME algorithm with $(\mathbf{X},\mathbf{V})$ needs fewer computations at each iteration.. Thus, we rewrite the snapshots or measurements $\mathbf{Y}$ in (2) by [2, 4, 3]
@@
where $\mathbf{x}(t)\sim\mathcal{CN}(\mathbf{0},\mathbf{I}_{M})$ is composed of $M$ common factors and uncorrelated with $\mathbf{v}(t)$. Then, the complete-data LLF is
$l(\mathbf{X},\mathbf{V};\mathbf{Q})$ $=$ $\ln p(\mathbf{X})+\ln p(\mathbf{V};\mathbf{Q})$
@@
where $c$ is a term unrelated to any free parameters in $(\mathbf{S},\mathbf{Q})$, $\hat{\mathbf{R}}_{v}=(1/L)\sum_{t=1}^{L}\mathbf{v}(t)\mathbf{v}^{H}(t)$. With (4.1), we can construct the ECME algorithm, whose E-step and two CM-steps at the $k$th iteration are derived below.
@@ 4.1.1 First CM-step:
Given $\mathbf{Q}=\mathbf{Q}^{(k-1)}$, let this step only obtain $\mathbf{S}^{(k)}$ by minimizing the actual or negative LLF $f(\mathbf{S},\mathbf{Q})$ with respect to $\mathbf{S}$, i.e.,
@@
This problem is classic [2, 4, 3] but the classic solutions in [18, 19] may not exist. Thus, we resort to the FAAN method by
@@ 4.1.2 E-step:
Given $\mathbf{Y}$, $\mathbf{S}^{(k)}$, and $\mathbf{Q}^{(k-1)}$, this step finds the expectation of $l(\mathbf{X},\mathbf{V};\mathbf{Q})$ in (4.1) by
$l(\mathbf{Q};\boldsymbol{\Omega}^{(k)})$ $=$ $\mathrm{E}\big{\{}l(\mathbf{X},\mathbf{V};\mathbf{Q})\mid\mathbf{Y};\boldsymbol{\Omega}^{(k)}\}$
@@
where $\boldsymbol{\Omega}^{(k)}=(\mathbf{S}^{(k)},\mathbf{Q}^{(k-1)})$, $\mathrm{E}\{\cdot\}$ and $\mathrm{D}\{\cdot\}$ means expectation and covariance, respectively. Moreover,
$\hat{\mathbf{R}}^{(k)}_{v}=\mathrm{E}\big{\{}\hat{\mathbf{R}}_{v}\mid\mathbf{Y};\boldsymbol{\Omega}^{(k)}\big{\}}=[\boldsymbol{\delta}^{(k)}]^{H}\hat{\mathbf{R}}\boldsymbol{\delta}^{(k)}+\boldsymbol{\Delta}^{(k)}\succ\mathbf{0}_{N}$
where $\mathbf{C}^{(k)}=\mathbf{S}^{(k)}[\mathbf{S}^{(k)}]^{H}+\mathbf{Q}^{(k-1)}$, $\boldsymbol{\delta}^{(k)}=[\mathbf{C}^{(k)}]^{-1}\mathbf{Q}^{(k-1)}$ [^4] We do not adopt $\mathbf{C}^{(k)}]^{-1}=\mathbf{q}^{(k)}-\mathbf{q}^{(k)}\mathbf{S}^{(k)}(\boldsymbol{\Lambda}^{(k)}+\mathbf{I}_{M})^{-1}[\mathbf{S}^{(k)}]^{H}\mathbf{q}^{(k)}$ with $\mathbf{q}^{(k)}=[\mathbf{Q}^{(k-1)}]^{-1}$ since numerical results show its low accuracy., and $\boldsymbol{\Delta}^{(k)}=\mathbf{Q}^{(k-1)}-\mathbf{Q}^{(k-1)}\boldsymbol{\delta}^{(k)}\succeq\mathbf{0}_{N}$. The conditional distribution of $\mathbf{v}(t)$ can be derived from [20], and
$\mathrm{E}\big{\{}\mathbf{v}(t)\mid\mathbf{Y};\boldsymbol{\Omega}^{(k)}\big{\}}$ $=$ $[\boldsymbol{\delta}^{(k)}]^{H}\mathbf{y}(t),$
$\mathrm{D}\big{\{}\mathbf{v}(t)\mid\mathbf{Y};\boldsymbol{\Omega}^{(k)}\big{\}}$ $=$ $\boldsymbol{\Delta}^{(k)}\succeq\mathbf{0}_{N}.$
@@ 4.1.3 Second CM-step:
This step only obtains $\mathbf{Q}^{(k)}$ by maximizing $l(\mathbf{Q};\boldsymbol{\Omega}^{(k)})$ in (4.1.2) with respect to $\mathbf{Q}$, i.e.,
@@
$\Big{(}\min_{\sigma_{n}^{2}>0}\ln(\sigma_{n}^{2})+[\hat{\mathbf{R}}_{v}^{(k)}]_{n,n}/\sigma^{2}_{n}\Big{)}~~n=1,\dots,N.$ (18)
The solutions are
@@
where $\mathrm{diag}(\cdot)$ is the matrix composed of the diagonal part. The details of the ECME algorithm are in Algorithm 2.
@@ 4.2 Convergence and Complexity
From [7, 21], we know that the ECME algorithm satisfies certain regularity conditions and its sequence $\big{\{}\big{(}\mathbf{S}^{(k)},\mathbf{Q}^{(k)}\big{)}\big{\}}$ always converges to a stationary point of $f(\mathbf{S},\mathbf{Q})$. Fortunately, simulation results in [10] and the subsequent section show that although the objective function $f(\mathbf{S},\mathbf{Q)}$ in problem (5) is non-convex, it seems to have a favourable landscape, which makes its minimization by the FAAN method or the ECME algorithm a relatively straightforward task. Moreover, Eq. (15) and Eq. (4.1.3) guarantee that the solution provided by the ECME algorithm is “proper" for problem (5) [18, 22].
By comparing Algorithm 1 and Algorithm 2, we can find that at the $k$th iteration, the computational complexity of Algorithm 1 lies in the eigen-decomposition of $\tilde{\mathbf{R}}^{(k)}\in\mathbb{C}^{N\times N}$ in (9) and the inverse of $\big{(}\mathbf{U}^{(k)}\boldsymbol{\Lambda}^{(k)}[\mathbf{U}^{(k)}]^{H}+\mathbf{I}_{N}\big{)}\in\mathbb{C}^{N\times N}$ in (3.2). Also, the computational complexity of Algorithm 2 lies in the eigen-decomposition of $\tilde{\mathbf{R}}^{(k)}\in\mathbb{C}^{N\times N}$ in (15) and the inverse of $\mathbf{C}^{(k)}\in\mathbb{C}^{N\times N}$ in (4.1.3). Thus, the ECME algorithm has almost the same computational complexity at each iteration as the FAAN method.
@@ 5 Numerical Results
Simulation results are used to show and compare Algorithm 1 and Algorithm 2. Let $N=6$, $M=2$, $\mathbf{P}=10\times\mathbf{I}_{M}$, $\mathbf{Q}=\mathrm{diag}\{10,2,3,2,1,3\}$, $\theta_{1}=60^{\circ}$, $\theta_{2}=120^{\circ}$, and $K_{2}=100$. Moreover, $\mathbf{Q}^{(0)}=\mathbf{I}_{N}$.
Without loss of generality, Fig. 1 illustrates the convergence performances of both algorithms given the same snapshots. Here, $L=100$. We can clearly see that the ECME algorithm yields faster convergence than the FAAN method. Because the ECME algorithm has almost the same computational complexity at each iteration as the FAAN method, it is computationally more efficient.
Figure 1: Convergence comparison of the FAAN method and the ECME algorithm.

### v3

@@ 4.1 Algorithm Procedure
For problem (6), we do not adopt the classic complete data $(\mathbf{Y},\mathbf{X})$ with $\mathbf{X}=\big[\mathbf{x}(1)~\cdots~\mathbf{x}(L)\big]$ in [rubin, liu2] and instead introduce the new complete data $(\mathbf{X},\mathbf{V})$ with $\mathbf{V}=\big[\mathbf{v}(1)~\cdots~\mathbf{v}(L)\big]$ [^3] The complete-data LLF of $(\mathbf{Y},\mathbf{X})$ is $l(\mathbf{X};\mathbf{S},\mathbf{Q})=\ln p(\mathbf{Y},\mathbf{X};\mathbf{S},\mathbf{Q})$ $={\sum}_{t=1}^{L}\Big[\ln p\big(\mathbf{y}(t)\mid\mathbf{x}(t);\mathbf{S},\mathbf{Q}\big)+\ln p\big(\mathbf{x}(t)\big)\Big]$ $=c-L\ln\mathrm{Det}(\mathbf{Q})-{\sum}_{t=1}^{L}\big[\mathbf{y}(t)-\mathbf{Sx}(t)\big]^{H}\mathbf{Q}^{-1}\big[\mathbf{y}(t)-\mathbf{Sx}(t)\big]$ where $c$ is a term unrelated to any free parameters in $(\mathbf{S},\mathbf{Q})$, and more complex than that of $(\mathbf{X},\mathbf{V})$ in (14). Numerical results show that the ECME algorithm with $(\mathbf{X},\mathbf{V})$ yields the same convergence as that with $(\mathbf{Y},\mathbf{X})$, but the ECME algorithm with $(\mathbf{X},\mathbf{V})$ needs fewer computations at each iteration.. Toward this end, we rewrite the snapshots or measurements $\mathbf{Y}$ in (2) by [lawley, basilevsky, mulaik]
@@
where $\mathbf{x}(t)\sim\mathcal{CN}(\mathbf{0},\mathbf{I}_{M})$ is composed of $M$ common factors and uncorrelated with $\mathbf{v}(t)$. Here, the statistical design of $\mathbf{x}(t)$ is based on $\mathbf{y}(t)\sim\mathcal{CN}(\mathbf{0},\mathbf{SS}^{H}+\mathbf{Q})$. Then, the complete-data LLF is
$l(\mathbf{X},\mathbf{V};\mathbf{Q})$ $=$ $\ln p(\mathbf{X},\mathbf{V};\mathbf{Q})=\ln p(\mathbf{X})+\ln p(\mathbf{V};\mathbf{Q})$ (14)
@@
where $c$ is a term unrelated to any free parameters in $(\mathbf{S},\mathbf{Q})$ and $\hat{\mathbf{R}}_{v}=(1/L)\sum_{t=1}^{L}\mathbf{v}(t)\mathbf{v}^{H}(t)$. With (14), we can construct the ECME algorithm, whose E-step and two CM-steps at the $k$th iteration are derived below.
@@ 4.1.1 First CM-step:
Given $\mathbf{Q}=\mathbf{Q}^{(k-1)}\succ\mathbf{0}_{N}$, let this step only obtain $\mathbf{S}^{(k)}$ by directly minimizing the objective function $f(\mathbf{S},\mathbf{Q})$ in problem (6) with respect to $\mathbf{S}$, i.e.,
@@
This problem is classic [lawley, basilevsky, mulaik] but the classic solutions in [joreskog, bartholomew] may not exist. Thus, we resort to the FAAN method by
@@ 4.1.2 E-step:
Given $\mathbf{Y}$, $\mathbf{S}^{(k)}$, and $\mathbf{Q}^{(k-1)}$, this step finds the conditional expectation of the complete-data LLF $l(\mathbf{X},\mathbf{V};\mathbf{Q})$ in (14) by
$l(\mathbf{Q};\boldsymbol{\Omega}^{(k)})=\mathrm{E}\big\{l(\mathbf{X},\mathbf{V};\mathbf{Q})\mid\mathbf{Y};\boldsymbol{\Omega}^{(k)}\}$ (17)
@@
where $\boldsymbol{\Omega}^{(k)}=(\mathbf{S}^{(k)},\mathbf{Q}^{(k-1)})$, $\mathrm{E}\{\cdot\}$ and $\mathrm{D}\{\cdot\}$ mean the expectation and covariance operators, respectively. Moreover,
$\hat{\mathbf{R}}^{(k)}_{v}$ $=$ $\mathrm{E}\big\{\hat{\mathbf{R}}_{v}\mid\mathbf{Y};\boldsymbol{\Omega}^{(k)}\big\}$ (18)
$=$ $\frac{1}{L}\sum_{t=1}^{L}\mathrm{E}\big\{\mathbf{v}(t)\mathbf{v}^{H}(t)\mid\mathbf{Y};\boldsymbol{\Omega}^{(k)}\big\}$
$=$ $\frac{1}{L}\sum_{t=1}^{L}\Big(\boldsymbol{\Sigma}^{(k)}+\boldsymbol{\mu}^{(k)}(t)\big[\boldsymbol{\mu}^{(k)}(t)\big]^{H}\Big)$
$=$ $\boldsymbol{\Sigma}^{(k)}+\frac{1}{L}\sum_{t=1}^{L}\boldsymbol{\mu}^{(k)}(t)[\boldsymbol{\mu}^{(k)}(t)]^{H}$
$=$ $\boldsymbol{\Delta}^{(k)}+[\boldsymbol{\delta}^{(k)}]^{H}\hat{\mathbf{R}}\boldsymbol{\delta}^{(k)}\succ\mathbf{0}_{N}$
where $\mathbf{C}^{(k)}=\mathbf{S}^{(k)}[\mathbf{S}^{(k)}]^{H}+\mathbf{Q}^{(k-1)}\succ\mathbf{0}_{N}$, $\boldsymbol{\delta}^{(k)}=[\mathbf{C}^{(k)}]^{-1}\mathbf{Q}^{(k-1)}$ [^4] We do not adopt $\mathbf{C}^{(k)}]^{-1}=\mathbf{q}^{(k)}-\mathbf{q}^{(k)}\mathbf{S}^{(k)}(\boldsymbol{\Lambda}^{(k)}+\mathbf{I}_{M})^{-1}[\mathbf{S}^{(k)}]^{H}\mathbf{q}^{(k)}$ with $\mathbf{q}^{(k)}=[\mathbf{Q}^{(k-1)}]^{-1}$ since numerical results show its low accuracy., and $\boldsymbol{\Delta}^{(k)}=\mathbf{Q}^{(k-1)}-\mathbf{Q}^{(k-1)}\boldsymbol{\delta}^{(k)}\succeq\mathbf{0}_{N}$. Additionally, the conditional distribution of $\mathbf{v}(t)$ can be derived from [rhodes], and
$\boldsymbol{\mu}^{(k)}(t)$ $=$ $\mathrm{E}\big\{\mathbf{v}(t)\mid\mathbf{Y};\boldsymbol{\Omega}^{(k)}\big\}=[\boldsymbol{\delta}^{(k)}]^{H}\mathbf{y}(t),$ (19a)
$\boldsymbol{\Sigma}^{(k)}$ $=$ $\mathrm{D}\big\{\mathbf{v}(t)\mid\mathbf{Y};\boldsymbol{\Omega}^{(k)}\big\}=\boldsymbol{\Delta}^{(k)}\succeq\mathbf{0}_{N}.$ (19b)
@@ 4.1.3 Second CM-step:
This step only obtains $\mathbf{Q}^{(k)}$ by maximizing the expected complete-data LLF $l(\mathbf{Q};\boldsymbol{\Omega}^{(k)})$ in (17) with respect to $\mathbf{Q}$, i.e.,
@@
$\min_{\sigma_{n}^{2}>0}\ln(\sigma_{n}^{2})+[\hat{\mathbf{R}}_{v}^{(k)}]_{n,n}/\sigma^{2}_{n},n=1,\dots,N.$ (21)
Let $f(\sigma^{2}_{n})=\ln(\sigma^{2}_{n})+[\hat{\mathbf{R}}_{v}^{(k)}]_{n,n}/\sigma^{2}_{n}$ with $\hat{\mathbf{R}}_{v}^{(k)}]_{n,n}>0$ and its derivative is
$f^{\prime}(\sigma^{2}_{n})$ $=$ $\frac{1}{\sigma^{2}_{n}}-\frac{[\hat{\mathbf{R}}_{v}^{(k)}]_{n,n}}{\sigma^{4}_{n}}=\frac{\sigma^{2}_{n}-[\hat{\mathbf{R}}_{v}^{(k)}]_{n,n}}{\sigma^{4}_{n}}$ (25)
[…]
which indicates that when $\sigma^{2}_{n}=[\hat{\mathbf{R}}_{v}^{(k)}]_{n,n}$, $f(\sigma^{2}_{n})$ reaches the minimum. Thus, the solutions in subproblems (21) are
@@
where $\mathrm{diag}(\hat{\mathbf{R}}_{v}^{(k)})$ is the diagonal matrix composed of the diagonal part of $\hat{\mathbf{R}}_{v}^{(k)}$. The details of the ECME algorithm are in Algorithm 2.
@@ 4.2 Convergence and Complexity
From [liu, wu, McLachlan], we know that since only the complex Gaussian distribution, whose PDF has a regular exponential-family form, is involved, the ECME algorithm satisfies the “space-filling” and other regularity conditions and its sequence $\big\{\big(\mathbf{S}^{(k)},\mathbf{Q}^{(k)}\big)\big\}$ always converges to a stationary point of $f(\mathbf{S},\mathbf{Q})$. Theoretically, $f(\mathbf{S},\mathbf{Q})$ has multiple stationary points, so an accurate $\mathbf{Q}^{(0)}$ should be required by the ECME algorithm. Fortunately, simulation results in [stoica] and the subsequent section show that although the objective function $f(\mathbf{S},\mathbf{Q)}$ in problem (6) is non-convex, it seems to have a favourable landscape under moderately nonuniform noise, which makes its minimization by the FAAN method or the ECME algorithm a relatively straightforward task. However, since the second step at each iteration of the FAAN method sequentially (one by one) updates the estimates of $\sigma_{1},\dots,\sigma_{N}$ along the fixed coordinate directions, its sequence $\big\{\big(\mathbf{S}^{(k)},\mathbf{Q}^{(k)}\big)\big\}$ gets stuck at an undesired limit point more easily when the noise is immoderately nonuniform, which will be shown in Fig. 3.
We utilize the following proposition to guarantee that the solution provided by the ECME algorithm is “proper” for problem (6) [joreskog, adachi].
Proposition 1.
In the ECME algorithm, $\mathbf{S}^{(k)}[\mathbf{S}^{(k)}]^{H}\succeq\mathbf{0}_{N}$ and $\mathbf{Q}^{(k)}\succ\mathbf{0}_{N}$ for arbitrary $k\geq 1$ if $\hat{\mathbf{R}}\succ\mathbf{0}_{N}$ and $\mathbf{Q}^{(0)}\succ\mathbf{0}_{N}$.
Proof.
We utilize the mathematical induction method. When $\hat{\mathbf{R}}\succ\mathbf{0}_{N}$ and $\mathbf{Q}^{(k-1)}\succ\mathbf{0}_{N}$, Eq. (16) directly indicates
$\mathbf{S}^{(k)}[\mathbf{S}^{(k)}]^{H}$
$=[\mathbf{Q}^{(k-1)}]^{1/2}\mathbf{U}^{(k)}\boldsymbol{\Lambda}^{(k)}[\mathbf{U}^{(k)}]^{H}[\mathbf{Q}^{(k-1)}]^{1/2}\succeq\mathbf{0}_{N}$
and then $\mathbf{C}^{(k)}=\mathbf{S}^{(k)}[\mathbf{S}^{(k)}]^{H}+\mathbf{Q}^{(k-1)}\succ\mathbf{0}_{N}$ and $\boldsymbol{\delta}^{(k)}=[\mathbf{C}^{(k)}]^{-1}\mathbf{Q}^{(k-1)}$ are of full rank. As a result, $\boldsymbol{\delta}^{(k)}]^{H}\hat{\mathbf{R}}\boldsymbol{\delta}^{(k)}\succ\mathbf{0}_{N}\Rightarrow\hat{\mathbf{R}}^{(k)}_{v}=\boldsymbol{\Delta}^{(k)}+[\boldsymbol{\delta}^{(k)}]^{H}\hat{\mathbf{R}}\boldsymbol{\delta}^{(k)}\succ\mathbf{0}_{N}$ in (18) due to the covariance matrix $\boldsymbol{\Delta}^{(k)}\succeq\mathbf{0}_{N}$. Finally, we obtain $\mathbf{Q}^{(k)}=\mathrm{diag}\big(\hat{\mathbf{R}}^{(k)}_{v}\big)\succ\mathbf{0}_{N}$ in (4.1.3). The proof is completed. ∎
By comparing Algorithm 1 and Algorithm 2, Table 2 and Table 2 give their main operations of the $k$th iteration. From both tables, we can know that the computational complexities of each iteration in both algorithms involve $\mathit{\mathcal{O}}(N^{3})$ when $N$ is large [Izadkhah]. Thus, the ECME algorithm has almost the same computational complexity at each iteration as the FAAN method.
Table 1: Computational complexity of the $k$th iteration in the FAAN method
Main Operation eigen-decomposition of $\tilde{\mathbf{R}}^{(k)}$ in (10) $\boldsymbol{\Gamma}^{(k)}=\big(\mathbf{U}^{(k)}\boldsymbol{\Lambda}^{(k)}[\mathbf{U}^{(k)}]^{H}+\mathbf{I}_{N}\big)^{-1}$ in (12) the $b_{n}^{(k)}$’s in (12)
Complexity $\mathit{\mathcal{O}}(N^{3})$ $\mathit{\mathcal{O}}(N^{3})$ $\mathit{\mathcal{O}}(K_{2}N^{2})$
Table 2: Computational complexity of the $k$th iteration in the ECME algorithm
Main Operation eigen-decomposition of $\tilde{\mathbf{R}}^{(k)}$ in (16) $\mathbf{C}^{(k)}]^{-1}$ in (18) matrix multiplications in (18)
Complexity $\mathit{\mathcal{O}}(N^{3})$ $\mathit{\mathcal{O}}(N^{3})$ $\mathit{\mathcal{O}}(N^{3})$
@@ 5 Numerical Results
Simulation results are used to show and compare Algorithm 1 and Algorithm 2. Unless otherwise specified in the following figures, let $N=6$, $M=2$, $\theta_{1}=60^{\circ}$, $\theta_{2}=120^{\circ}$, $\mathbf{P}=10\times\mathbf{I}_{M}$, $\mathbf{Q}=\mathrm{diag}\{10,2,3,2,1,3\}$, $K_{2}=100$, $L=100$, $K=K_{1}=100$, and $\mathbf{Q}^{(0)}=\mathbf{I}_{N}$.
5.1 Convergence Evaluation
Without loss of generality, Fig. 1 illustrates the convergence performances of both algorithms given the same snapshots. Here, the upper subfigure is with $\theta_{1}=90^{\circ}$ and $\theta_{2}=100^{\circ}$. We can clearly see that the ECME algorithm yields faster convergence than the FAAN method. Because the ECME algorithm has almost the same computational complexity at each iteration as the FAAN method, it is computationally more efficient.
Figure 1: Convergence comparison of the FAAN method and the ECME algorithm
@@ 5.1 Convergence Evaluation
Fig. 2 and Fig. 3 show two scatter plots of $(\hat{\theta}_{1},\hat{\theta}_{2})$’s from both algorithms under $100$ independent realizations. The same snapshots of each realization are processed by both algorithms. In Fig. 2, the noise is moderately nonuniform and all the $(\hat{\theta}_{1},\hat{\theta}_{2})$’s are centered on the true value $(60^{\circ},120^{\circ})$ since for each realization, both sequences $\{\mathbf{S}^{(k)}\mathbf{S}^{(k)}\}$’s from the FAAN method and the ECME algorithm converge to the solution of $\mathbf{S}\mathbf{S}^{H}$ in problem (6). However, we know that under the same number of iterations $K=K_{1}$, both algorithms provide different $\hat{\mathbf{S}}\hat{\mathbf{S}}^{H}$’s for each realization, and the $\hat{\mathbf{S}}\hat{\mathbf{S}}^{H}$ provided by the ECME algorithm is closer to the solution of $\mathbf{S}\mathbf{S}^{H}$ in problem (6). In Fig. 3, $\mathbf{Q}=\mathrm{diag}\{10,2,3000,2,1,3\}$, the numbers of desired points from the FAAN method and the ECME algorithm are $53$ and $98$, respectively. This implies that when the noise is immoderately nonuniform, the convergence of the FAAN method to an undesired limit point is easier. That is to say, the convergence of the ECME algorithm to the global optimum is easier.
Figure 2: Scatter plot of $(\hat{\theta}_{1},\hat{\theta}_{2})$’s from both algorithms in moderately nonuniform noise
Figure 3: Scatter plot of $(\hat{\theta}_{1},\hat{\theta}_{2})$’s from both algorithms in immoderately nonuniform noise
When $\mathbf{P}=\begin{bmatrix}10~10\\ 10~10\end{bmatrix}$, i.e., the source signals are coherent, Fig. 4 shows a scatter plot of $(\hat{\theta}_{1},\hat{\theta}_{2})$’s from both algorithms under $100$ independent realizations. The same snapshots of each realization are processed by both algorithms. In this figure, many $(\hat{\theta}_{1},\hat{\theta}_{2})$’s from both algorithms are not centered on the true value $(60^{\circ},120^{\circ})$ although the noise is moderately nonuniform. This indicates that in the presence of coherent source signals, the classic MLFA problem (6) cannot be employed directly to estimate the signal subspace and the spatial smoothing technique should be considered [Shan, Pillai].
Figure 4: Scatter plot of $(\hat{\theta}_{1},\hat{\theta}_{2})$’s from both algorithms with coherent source signals
5.2 Accuracy Evaluation
In this subsection, the root mean square error (RMSE) performances of DOA estimation related to both algorithms are compared with the Cramer–Rao lower bound (CRLB) [pesavento, gershman]. To guarantee the convergence of both algorithms to the global optimum, let the source signals be noncoherent and the noise be moderately nonuniform. In the following figures, each RMSE is based on $1000$ independent realizations and the RMSE of $\hat{\theta}_{m}$ is
$\mathrm{RMSE}(\hat{\theta}_{m})$ $=$ $\sqrt{\frac{1}{1000}\sum_{i=1}^{1000}\Big(\hat{\theta}_{m}^{i}-\theta_{m}\Big)^{2}}$
$=$ $10\lg\Big[\frac{1}{1000}\sum_{i=1}^{1000}\Big(\hat{\theta}_{m}^{i}-\theta_{m}\Big)^{2}\Big]~\mathrm{(dB)},$
where $m=1,\dots,M$ and $\hat{\theta}_{m}^{i}$ denotes $\hat{\theta}_{m}$ obtained from the $i$th realization.